OpenStudy (anonymous):

tanx +square root 3= sec x

OpenStudy (lgbasallote):

isolate the x first... tanx - secx = - sqrt 3 \[\large \frac {sinx}{cosx} - \frac {1}{cosx} = -\sqrt{3}\] \[\large \frac {sinx-1}{cosx} = -\sqrt{3}\] hmmmm..i may be on the right track but im stuck :/

OpenStudy (lgbasallote):

\[\large \frac {sinx - (\sin ^{2}x + \cos ^{2}x)}{cosx} = -\sqrt{3}\]

OpenStudy (lgbasallote):

i think im getting somewhere...

OpenStudy (callisto):

tanx +square root 3= sec x tanx = secx - sqrt 3 Square both sides (tanx)^2 = (secx - sqrt 3)^2 Given that (tanx)^2 = (secx)^2 - 1 (secx)^2 - 1 = (secx)^2 - 2sqrt 3 (secx) +3 0 = -2sqrt3 (secx) +4 secx = 2/sqrt 3 .....

OpenStudy (anonymous):

ok I have to find all solutions in between the interval 0<x<360 degrees for this problem

OpenStudy (lgbasallote):

ill leave @Callisto to this <tips hat>

OpenStudy (callisto):

Hold on, I need to check... sorry, just a minute

OpenStudy (anonymous):

no problem

OpenStudy (callisto):

I don't know why... when i sub my solution back, i can't get the answer...:(

OpenStudy (anonymous):

yea this is a troublesome problem for me too

OpenStudy (callisto):

secx = 2/sqrt 3 x = 30(rejected) or x =330 I found that another solution is 150.. but I don't know how to get it by solving the equation. Give me some more time..

OpenStudy (anonymous):

ok

OpenStudy (callisto):

Alright, I got it tanx +square root 3= sec x Square both sides (tanx+sqrt3)^2 = (secx)^2 Given that (tanx)^2 +1 = (secx)^2 (tanx)^2+2(sqrt3)(tanx) +(sqrt3)^2 = (tanx)^2 +1 2(sqrt3)(tanx) +3= 1 (2sqrt3) tanx = -2 tanx = -2 / (2sqrt3) tanx = -1/sqrt3 x = 150 or x=330

OpenStudy (anonymous):

lifesaver, thank you!!

OpenStudy (callisto):

Do you understand it?